OpenStudy (anonymous):

multiply (9+sqrt8)(9-sqrt8)

OpenStudy (anonymous):

73

OpenStudy (ashish56):

\[(9+\sqrt{8})(9-\sqrt{8})=9^{2}-(\sqrt{8})^{2}=81-8=73\]